# Distributed Thompson sampling under constrained communication

### v1

@@ C-A Theoretical Result
In the following theorem, we provide a bound for Bayesian Simple Regret for an $M$ agent system with communication graph $G.$ After proving Theorem 3.1, we will introduce and discuss its implication, Corollary 3.2, which is our main convergence result.
@@
Proof. The structure of our proof follows techniques from Kandasamy et al. [13]. We aim to provide a bound on Bayesian Simple Regret. Our proof begins by noting that we can bound Bayesian Simple Regret the average expected regret of any subset of agents. We then decompose this into three sums, each of which utilize a confidence function $U_{t,i}(\cdot).$ We bound each of these sums by incorporating notions from information theory, which allows us to use information gain constants introduced in Equations 3 and 2 to analyze the efficacy of the Gaussian process.
@@
Recall from equation 1 that $R_{SB}(t)=\min_{i\in\{1,2,\cdots,M\},\tau\in\{1,2,\cdots,t\}}\mathbb{E}[f_{i}(x^{*})-f_{i}(x_{\tau,i})].$ Define $U_{t,i}(x)=\mu_{D_{t,i}}(x)+\beta_{t}^{1/2}\sigma_{D_{t,i}}(x).$ Note that, since $R_{SB}(t)$ represents a minimum value of the expression $\mathbb{E}[f_{i}(x^{*})-f_{i}(x_{\tau,i})],$ it is upper bounded by the average of this expression across any subset of agents. Enumerate the agents in the complete graph $G_{s}=(V_{s},E_{s})$ as $\{1,2,\cdots,|V_{s}|\},$ and thus we reindex the $M$ agents such that agents $i\in\{1,2,\cdots,|V_{s}|\}\in V_{s}$ and agents $j\in\{|V_{s}|+1,\cdots,M\}\in V\backslash V_{s}.$ Thus we can write that
$R_{SB}(t)\leq\frac{1}{t|V_{s}|}\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[f_{i}(x^{*})-f_{i}(x_{\tau,i})].$
We will then focus on upper bounding the expression $\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[f_{i}(x^{*})-f_{i}(x_{\tau,i})].$
$\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[f_{i}(x^{*})-f_{i}(x_{\tau,i})]$
$=\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[f_{i}(x^{*})-U_{\tau,i}(x^{*})+U_{\tau,i}(x^{*})-U_{\tau,i}(x_{\tau,i})$
$\hskip 56.9055pt+U_{\tau,i}(x_{\tau,i})-f_{i}(x_{\tau,i})]$
$=\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\underbrace{\mathbb{E}[f_{i}(x^{*})-U_{\tau,i}(x^{*})]}_{S1}+\underbrace{\mathbb{E}[U_{\tau,i}(x^{*})-U_{\tau,i}(x_{\tau,i})]}_{S2}$
$\hskip 56.9055pt+\underbrace{\mathbb{E}[U_{\tau,i}(x_{\tau,i})-f_{i}(x_{\tau,i})]}_{S3}$
We will now bound each of these sums.
S1. Let’s begin by upper bounding the sum                       $S1=\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[f_{i}(x^{*})-U_{\tau,i}(x^{*})].$
$S1$ $=\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[f_{i}(x^{*})-U_{\tau,i}(x^{*})]$ (4)
$\leq\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[\mathbb{I}\{f_{i}(x^{*})>U_{\tau,i}(x^{*})\}\left(f_{i}(x^{*})-U_{\tau,i}(x^{*})\right)]$ (5)
$\leq\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\sum_{x\in\mathcal{X}}\frac{e^{-\beta_{\tau}/2}}{\sqrt{2\pi}}$ (6)
$\leq\sum_{\tau=1}^{t}\frac{1}{\tau^{2}\sqrt{2\pi}}\leq\frac{\sqrt{2}\pi^{3/2}}{12}$ (7)
Line (5) upper bounds $S1$ by positive terms, Line (6) utilizes Lemma A.1 in Appendix A to bound the expectation of positive terms in a normal distribution. Line (7) results from substituting for $\beta_{\tau},$ and the fact that $\sum_{j=1}^{\infty}\frac{1}{j^{2}}=\frac{\pi^{2}}{6}.$ Therefore, we have established an upper bound for $S1$.
S2. Let’s evaluate the expression
$S2=\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[U_{\tau,i}(x^{*})-U_{\tau,i}(x_{\tau,i})].$ First let’s focus on the interior of the summation, $\mathbb{E}[U_{\tau,i}(x^{*})-U_{\tau,i}(x_{\tau,i})].$ We will proceed to show that this expression evaluates to $0$
S3. Lastly, we bound
$S3=\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[U_{\tau,i}(x_{\tau,i})\!-\!f_{i}(x_{\tau,i})].$ For the evaluation of this sum, we will introduce some additional notation. Denote $\bar{D}_{t}=\{(x_{\tau,j},y_{\tau,j})\}_{j\in V_{s},\tau<t}.$ We can think of $\bar{D}_{t}$ as representing the data acquired by round $t$ in a batch setting exclusively by agents in $V_{s}$. Additionally, define $\sigma_{t,i}(x):=\sigma(x)\mid\bar{D}_{t}\cup\left\{(x_{t,j},y_{t,j})\right\}_{j\in V_{s},j<i}.$ We pick to define $\sigma_{t,i}(\cdot)$ in this way to impose an ordering of the acquired data; in this sense, $\sigma_{t,i}(\cdot)$ depends on $(t-1)|V_{s}|+i-1$ previously sampled points. With this additional notation, we are well-equipped to upper bound $S3.$
$S3$ $=\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[U_{\tau,i}(x_{\tau,i})-f_{i}(x_{\tau,i})]$ (8)
$=\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[\mathbb{E}[\mu_{D_{\tau,i}}(x_{\tau,i})+\beta_{\tau}^{1/2}\sigma_{D_{\tau,i}}(x_{\tau,i})$
$\hskip 49.79231pt-f_{i}(x_{\tau,i})]|D_{\tau,i}]]$ (9)
$=\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[\beta_{\tau}^{1/2}\sigma_{D_{\tau,i}}(x_{\tau,i})]$ (10)
$\leq\beta_{t}^{1/2}\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[\sigma_{D_{\tau,i}}(x_{\tau,i})]$ (11)
$\leq\beta_{t}^{1/2}\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[\sigma_{\bar{D}_{\tau}}(x_{\tau,i})]$ (12)
$\leq\beta_{t}^{1/2}\mathbb{E}\left[\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\sigma_{\tau,i}(x_{\tau,i})\exp(I(f;\{y_{\tau,j}\}_{j<i}|y_{\bar{D}_{\tau}})\right]$ (13)
$\leq\beta_{t}^{1/2}\mathbb{E}\left[\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\sigma_{\tau,i}(x_{\tau,i})\xi_{i}^{1/2}\right]$ (14)
$\leq\beta_{t}^{1/2}\xi_{|V_{s}|}^{1/2}\mathbb{E}\left[\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\sigma_{\tau,i}(x_{\tau,i})\right]$ (15)
$\leq\beta_{t}^{1/2}\xi_{|V_{s}|}^{1/2}\mathbb{E}\left[\Big{(}t|V_{s}|\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\sigma^{2}_{\tau,i}(x_{\tau,i})\Big{)}^{1/2}\right]$ (16)
$\leq\beta_{t}^{1/2}\xi_{|V_{s}|}^{1/2}\sqrt{\frac{2t|V_{s}|\Psi_{t|V_{s}|}}{\log(1+\sigma_{\epsilon}^{-2})}}$ (17)
$\leq\sqrt{\frac{2\xi_{|V_{s}|}t|V_{s}|\beta_{t}\Psi_{t|V_{s}|}}{\log(1+\sigma_{\epsilon}^{-2})}}$ (18)
Lines (10) and (11) follow by the law of total expectation. Line (12) follows by noting that $\beta_{\tau}$ is increasing with $\tau.$ Recall that because $G_{s}$ is a subgraph of $G,$ for all agents $i\in V_{s},$ $V_{s}\subset(N(i)\cup i).$ Therefore, $\bar{D}_{\tau}$ is contained in $D_{\tau,i}$ for all agents $i\in V_{s}.$ In essence, $\bar{D}_{\tau}\subset D_{\tau,i}\implies\sigma_{\bar{D}_{\tau}}\geq\sigma_{D_{\tau,i}}.$ Line (13) applies this property. Line (14) follows from Lemma A.4 in Appendix A, which introduces $\xi_{i}$ to bound the information gain of the current set of evaluations. Line (15) follows from the definition of $\xi_{i},$ which was stated in Equation 3. Line (16) is a consequence of the fact that $\xi_{i}$ is increasing with $i,$ and Line (17) follows from application of the Cauchy-Schwarz inequality. Finally, Line (18) is a consequence of Lemma A.3, which bounds the sums of the posterior variances by the MIG term.
Figure 1: Simulations of regret analysis with varying agent quantity on star graph configuration, on Ackley (left) and Rosenbrock (right) objective functions.
Recall that $\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[f_{i}(x^{*})-f_{i}(x_{\tau,i})]=S1+S2+S3.$ Thus, we have shown that $\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[f_{i}(x^{*})-f_{i}(x_{\tau,i})]\leq\frac{\sqrt{2}\pi^{3/2}}{12}+\sqrt{\frac{2\xi_{|V_{s}|}t|V_{s}|\beta_{t}\Psi_{t|V_{s}|}}{\log(1+\sigma_{\epsilon}^{-2})}}.$ However, we wish to bound $R_{SB}(t).$ Since $R_{SB}(t)\leq\frac{1}{t|V_{s}|}\sum_{\tau=1}^{t}\sum_{i=1}^{|V_{s}|}\mathbb{E}[f_{i}(x^{*})-f_{i}(x_{\tau,i})],$ it directly follows that $R_{SB}(t)\leq\frac{\sqrt{2}\pi^{3/2}}{12t|V_{s}|}+\sqrt{\frac{2\xi_{|V_{s}|}\beta_{t}\Psi_{t|V_{s}|}}{t|V_{s}|\log(1+\sigma_{\epsilon}^{-2})}},$ thus concluding our proof.
@@
Picking $G_{s}$ to be the largest complete subgraph of the communication network $G$ yields the following corollary.
@@
Suppose $k(x,x^{\prime})\leq 1$ for all $x,x^{\prime}.$ Let $G_{max}=(V_{max},E_{max})$ be the largest complete subgraph of $G.$ Then the Bayesian Simple Regret after $t$ timesteps satisfies $R_{SB}(t)\leq\frac{C_{1}}{t|V_{max}|}+\sqrt{\frac{C_{2}\xi_{|V_{max}|}\beta_{t}\Psi_{t|V_{max}|}}{t|V_{max}|}},$ where $\beta_{t}=2\log(t^{2}|V_{s}||\mathcal{X}|),$ $C_{1}=\frac{\sqrt{2}\pi^{3/2}}{12},$ and $C_{2}=\frac{2}{\log(1+\sigma_{\epsilon}^{-2})}.$
From the corollary, we see that $R_{SB}(t)=\tilde{O}\left(\sqrt{\frac{\xi_{|V_{max}|}\Psi_{t|V_{max}|}}{t|V_{max}|}}\right)$. We note that the term $\Psi_{t|V_{max}|}$ corresponds to the maximal mutual information gain from $t|V_{max}|$ observations, and that this quantity depends only logarithmically on $t|V_{max}|$ for standard kernels such as the squared exponential kernel. For more details, we refer to Appendix B. We note that the term $\xi_{|V_{max}|}$ is the price we pay for the delay in communication within the subgraph $G_{max}$, and is a standard term that arises in non-sequential Bayesian Optimization. By an appropriately designed initialization phase, this term can actually be reduced to $\tilde{O}(1)$ (see Appendix B.3 in [21] for more details). Thus, compared to the sequential case with $t$ rounds which has simple regret $\tilde{O}\left(\sqrt{\frac{1}{t}}\right)$ [9], our algorithm satisfies a regret of $\tilde{O}\left(\sqrt{\frac{1}{t|V_{max}|}}\right)$, i.e. an improvement of $\sqrt{\frac{1}{|V_{max}|}}$, demonstrating the benefit of the network structure.
D Numerical Experiments
D-A Simulation

### v3

@@ III-A Theoretical Result
We next provide a bound for Bayesian average regret for an $M$ agent system with communication graph $G.$
Theorem 3.1.
Suppose $k(x,x^{\prime})\leq 1$ for all $x,x^{\prime}.$ Let $\{G_{k}\}_{k\in\{1,\ldots,n\}}$ be a collection of $n$ disjoint complete subgraphs of communication graph $G=(V,E)$, where $G_{k}=(V_{k},E_{k}),$ and $\cup_{k\in\{1,\ldots,n\}}V_{k}=V.$ Then the Bayesian average regret after $t$ timesteps satisfies $R_{AB}(t)\leq\frac{1}{M}\sum_{k=1}^{n}|V_{k}|(\frac{C_{1}}{t|V_{k}|}+\sqrt{\frac{C_{2}\xi_{|V_{k}|}\beta_{t}\Psi_{t|V_{k}|}}{t|V_{k}|}}),$ where $\beta_{t}=2\log(t^{2}M|\mathcal{X}|),$ $C_{1}=\frac{\sqrt{2}\pi^{3/2}}{12},$ and $C_{2}=\frac{2}{\log(1+\sigma_{\epsilon}^{-2})}.$
Proof. The structure of our proof follows techniques from Kandasamy et al. [22]. We aim to provide a bound on Bayesian average regret. Our proof begins by noting that we can develop an expression for Bayesian average regret by considering the Bayesian average regret of specific subsets of agents. We then decompose this into three sums, each of which utilize a confidence function $U_{t,i}(\cdot).$ We bound each of these sums using notions from information theory, allowing us to use information gain constants introduced in Equations 3 and 4 to analyze the efficacy of the sampling process.
We bound the Bayesian average regret affiliated with agents in communication graph $G$ by bounding the Bayesian average regret within each complete subgraph of graph $G.$ Let $G=(V,E)$ be the communication graph for the $M$ agents. We can construct a collection of $n$ disjoint complete subgraphs $\{G_{k}\}_{k\in\{1,\ldots,n\}},$ where each $G_{k}=(V_{k},E_{k})$ is a subgraph of $G,$ with $\cup_{k\in\{1,2,\ldots,n\}}V_{k}=V.$
Recall from Equation 1 that Bayesian average regret $R_{AB}(t)=\frac{1}{tM}\sum_{\tau=1}^{t}\sum_{i=1}^{M}\mathbb{E}[f(x^{*})-f(x_{\tau,i})].$ We also introduce $R_{AB,k}(t)=\frac{1}{t|V_{k}|}\sum_{\tau=1}^{t}\sum_{i\in V_{k}}\mathbb{E}[f(x^{*})-f(x_{\tau,i})],$ which is the Bayesian average regret affiliated with agents in $V_{k}.$ Recalling the partition of the vertex set $V$ into $\{V_{k}\}_{k\in\{1,\ldots,n\}},$ we may rewrite Bayesian average regret as follows:
$R_{AB}(t)$ $=\frac{1}{tM}\sum_{\tau=1}^{t}\sum_{i=1}^{M}\mathbb{E}[f(x^{*})-f(x_{\tau,i})]$
$=\frac{1}{tM}\sum_{\tau=1}^{t}\sum_{k=1}^{n}\sum_{i\in V_{k}}\mathbb{E}[f(x^{*})-f(x_{\tau,i})]$
$=\frac{1}{tM}\sum_{k=1}^{n}\sum_{\tau=1}^{t}\sum_{i\in V_{k}}\mathbb{E}[f(x^{*})-f(x_{\tau,i})]$
$=\frac{1}{M}\sum_{k=1}^{n}|V_{k}|R_{AB,k}(t)$
Thus, it suffices to focus on bounding $R_{AB,k}(t).$ Define $U_{t,i}(x)=\mu_{D_{t,i}}(x)+\beta_{t}^{1/2}\sigma_{D_{t,i}}(x).$ To upper bound $R_{AB,k}(t),$ we can decompose the sum $\sum_{\tau=1}^{t}\sum_{i\in V_{k}}\mathbb{E}[f(x^{*})-f(x_{\tau,i})]$ as follows:
$\sum_{\tau=1}^{t}\sum_{i\in V_{k}}\mathbb{E}[f(x^{*})-f(x_{\tau,i})]$
$=\sum_{\tau=1}^{t}\sum_{i\in V_{k}}\mathbb{E}[f(x^{*})-U_{\tau,i}(x^{*})+U_{\tau,i}(x^{*})-U_{\tau,i}(x_{\tau,i})$
$\hskip 56.9055pt+U_{\tau,i}(x_{\tau,i})-f(x_{\tau,i})]$
$=\sum_{\tau=1}^{t}\sum_{i\in V_{k}}\underbrace{\mathbb{E}[f(x^{*})-U_{\tau,i}(x^{*})]}_{S1}+\underbrace{\mathbb{E}[U_{\tau,i}(x^{*})-U_{\tau,i}(x_{\tau,i})]}_{S2}$
$\hskip 56.9055pt+\underbrace{\mathbb{E}[U_{\tau,i}(x_{\tau,i})-f(x_{\tau,i})]}_{S3}$
We will now bound each of these sums.
S1. Let’s begin by upper bounding the sum                       $S1=\sum_{\tau=1}^{t}\sum_{i\in V_{k}}\mathbb{E}[f(x^{*})-U_{\tau,i}(x^{*})].$
$S1$ $=\sum_{\tau=1}^{t}\!\sum_{i\in V_{k}}\mathbb{E}[f(x^{*})-U_{\tau,i}(x^{*})]$ (5)
$\leq\sum_{\tau=1}^{t}\sum_{i\in V_{k}}\mathbb{E}\!\!\left[\mathbb{E}[\mathbb{I}\{f(x^{*})\!>\!U_{\tau,i}(x^{*})\}\!\!\left(f(x^{*})\!-\!U_{\tau,i}(x^{*})\right)\!\mid\!\mathcal{F}_{D_{t,i}}]\right]$ (6)
$\leq\sum_{\tau=1}^{t}\sum_{i\in V_{k}}\sum_{x\in\mathcal{X}}\frac{e^{-\beta_{\tau}/2}}{\sqrt{2\pi}}\leq\sum_{\tau=1}^{t}\frac{1}{\tau^{2}\sqrt{2\pi}}\leq\frac{\sqrt{2}\pi^{3/2}}{12}$ (7)
Line (6) upper bounds $S1$ by positive terms by making use of the indicator function $\mathbb{I}\{f(x^{*})>U_{\tau,i}(x^{*})\},$ which takes the value of $1$ when the condition $f(x^{*})>U_{\tau,i}(x^{*})$ is satisfied and 00 otherwise. Line (7) utilizes Lemma A.1 in Appendix A to bound the expectation of positive terms in a normal distribution. Additionally, Line (7) results from substituting for $\beta_{\tau},$ and the fact that $\sum_{j=1}^{\infty}\frac{1}{j^{2}}=\frac{\pi^{2}}{6}.$ Therefore, we have established an upper bound for $S1$.
S2. We evaluate the expression $S2=\sum_{\tau=1}^{t}\sum_{i\in V_{k}}\mathbb{E}[U_{\tau,i}(x^{*})-U_{\tau,i}(x_{\tau,i})].$ First let’s focus on the interior of the summation, $\mathbb{E}[U_{\tau,i}(x^{*})-U_{\tau,i}(x_{\tau,i})].$ We will proceed to show that this expression evaluates to $0$
S3. Lastly, we bound $S3=\sum_{\tau=1}^{t}\sum_{i\in V_{k}}\mathbb{E}[U_{\tau,i}(x_{\tau,i})\!-\!f(x_{\tau,i})].$ For the evaluation of this sum, we will introduce some additional notation. Denote $\bar{D}_{t}=\{(x_{\tau,j},y_{\tau,j})\}_{j\in V_{k},\tau<t}.$ We can think of $\bar{D}_{t}$ as representing the data acquired by round $t$ in a batch setting exclusively by agents in $V_{k}$. Additionally, define $\sigma_{t,i}(x):=\sigma(x)\mid\bar{D}_{t}\cup\left\{(x_{t,j},y_{t,j})\right\}_{j\in V_{k},j<i}.$ We pick to define $\sigma_{t,i}(\cdot)$ in this way to impose an ordering of the acquired data; in this sense, $\sigma_{t,i}(\cdot)$ depends on $(t-1)|V_{k}|+i-1$ previously sampled points. With this additional notation, we are well-equipped to upper bound $S3.$
$S3$ $=\sum_{\tau=1}^{t}\sum_{i\in V_{k}}\mathbb{E}[U_{\tau,i}(x_{\tau,i})-f(x_{\tau,i})]$ (8)
$=\sum_{\tau=1}^{t}\sum_{i\in V_{k}}\mathbb{E}[\mathbb{E}[\mu_{D_{\tau,i}}(x_{\tau,i})+\beta_{\tau}^{1/2}\sigma_{D_{\tau,i}}(x_{\tau,i})$
$\hskip 49.79231pt-f(x_{\tau,i})]|D_{\tau,i}]]$ (9)
$=\sum_{\tau=1}^{t}\sum_{i\in V_{k}}\mathbb{E}[\beta_{\tau}^{1/2}\sigma_{D_{\tau,i}}(x_{\tau,i})]$ (10)
$\leq\beta_{t}^{1/2}\sum_{\tau=1}^{t}\sum_{i\in V_{k}}\mathbb{E}[\sigma_{D_{\tau,i}}(x_{\tau,i})]$ (11)
$\leq\beta_{t}^{1/2}\sum_{\tau=1}^{t}\sum_{i\in V_{k}}\mathbb{E}[\sigma_{\bar{D}_{\tau}}(x_{\tau,i})]$ (12)
$\leq\beta_{t}^{1/2}\mathbb{E}\left[\sum_{\tau=1}^{t}\sum_{i\in V_{k}}\sigma_{\tau,i}(x_{\tau,i})\exp(I(f;\{y_{\tau,j}\}_{j<i}|y_{\bar{D}_{\tau}})\right]$ (13)
$\leq\beta_{t}^{1/2}\mathbb{E}\left[\sum_{\tau=1}^{t}\sum_{i\in V_{k}}\sigma_{\tau,i}(x_{\tau,i})\xi_{i}^{1/2}\right]$ (14)
$\leq\beta_{t}^{1/2}\xi_{|V_{k}|}^{1/2}\mathbb{E}\left[\sum_{\tau=1}^{t}\sum_{i\in V_{k}}\sigma_{\tau,i}(x_{\tau,i})\right]$ (15)
$\leq\beta_{t}^{1/2}\xi_{|V_{k}|}^{1/2}\mathbb{E}\left[\Big{(}t|V_{k}|\sum_{\tau=1}^{t}\sum_{i\in V_{k}}\sigma^{2}_{\tau,i}(x_{\tau,i})\Big{)}^{1/2}\right]$ (16)
$\leq\beta_{t}^{1/2}\xi_{|V_{k}|}^{1/2}\sqrt{\frac{2t|V_{k}|\Psi_{t|V_{k}|}}{\log(1+\sigma_{\epsilon}^{-2})}}\leq\sqrt{\frac{2\xi_{|V_{k}|}t|V_{k}|\beta_{t}\Psi_{t|V_{k}|}}{\log(1+\sigma_{\epsilon}^{-2})}}$ (17)
Lines (9) and (10) follow by the law of total expectation. Line (11) follows by noting that $\beta_{\tau}$ is increasing with $\tau.$ Recall that because $G_{k}$ is a subgraph of $G,$ for all agents $i\in V_{k},$ $V_{k}\subset\{N(i)\cup i\}.$ Therefore, $\bar{D}_{\tau}$ is contained in $D_{\tau,i}$ for all agents $i\in V_{k}.$ In essence, $\bar{D}_{\tau}\subset D_{\tau,i}\implies\sigma_{\bar{D}_{\tau}}\geq\sigma_{D_{\tau,i}}.$ Line (12) applies this property. Line (13) follows from Lemma A.4 in Appendix A, which introduces $\xi_{i}$ to bound the information gain of the current set of evaluations. Line (14) follows from the definition of $\xi_{i},$ which was stated in Equation 4. Line (15) is a consequence of the fact that $\xi_{i}$ is increasing with $i.$ Line (16) follows from application of the Cauchy-Schwarz inequality, and line (17) is a consequence of Lemma A.3, which bounds the sums of the posterior variances by the MIG term.
Recall that $\sum_{\tau=1}^{t}\sum_{i\in V_{k}}\mathbb{E}[f(x^{*})-f(x_{\tau,i})]=S1+S2+S3.$ Therefore, $R_{AB,k}(t)\leq\frac{\sqrt{2}\pi^{3/2}}{12t|V_{k}|}+\sqrt{\frac{2\xi_{|V_{k}|}\beta_{t}\Psi_{t|V_{k}|}}{t|V_{k}|\log(1+\sigma_{\epsilon}^{-2})}}.$
Equipped with a bound on $R_{AB,k}(t),$ we can revisit our expression for $R_{AB}(t).$
$R_{AB}(t)$ $=\frac{1}{M}\sum_{k=1}^{n}|V_{k}|R_{AB,k}(t)$
$\leq\frac{1}{M}\sum_{k=1}^{n}|V_{k}|\left(\frac{C_{1}}{t|V_{k}|}+\sqrt{\frac{C_{2}\xi_{|V_{k}|}\beta_{t}\Psi_{t|V_{k}|}}{t|V_{k}|}}\right),$
where $C_{1}=\frac{\sqrt{2}\pi^{3/2}}{12}$ and $C_{2}=\frac{2}{\log(1+\sigma_{\epsilon}^{-2})}.$ Thus, we have shown that $R_{AB}(t)\leq\frac{1}{M}\sum_{k=1}^{n}|V_{k}|\left(\frac{C_{1}}{t|V_{k}|}+\sqrt{\frac{C_{2}\xi_{|V_{k}|}\beta_{t}\Psi_{t|V_{k}|}}{t|V_{k}|}}\right),$ concluding our proof.
By picking $n$ to be the clique cover number of the graph $G$, Theorem 3.1 yields the following corollary.
Corollary 3.2.
Suppose $k(x,x^{\prime})\leq 1$ for all $x,x^{\prime}$. Let $\theta(G)$ and $\omega(G)$ denote the clique cover number and clique number of the graph $G$ respectively. Then, the Bayesian average regret after $t$ timesteps satisfies
$R_{AB}(t)\leq\frac{C_{1}\theta(G)}{Mt}+\frac{\sqrt{\theta(G)}\sqrt{C_{2}\xi_{\omega(G)}\beta_{t}\Psi_{t\omega(G)}}}{\sqrt{Mt}},$ where $\beta_{t},C_{1}$ and $C_{2}$ are as defined in Theorem 3.3.
Proof. The proof of Corollary 3.2 follows from (i) applying Cauchy-Schwarz to bound the term $\sum_{k=1}^{n}\sqrt{|V_{k}|}\leq\sqrt{n}\sqrt{\sum_{k=1}^{n}|V_{k}|}=\sqrt{Mn}$, (ii) picking $n$ to be the clique cover number of $G$, $\theta(G)$, and (iii) the fact that for any clique $G_{k}=(V_{k},E_{k})$ in $G$, $|V_{k}|\leq\omega(G)$, since $\omega(G)$ denotes the clique number of $G$ (i.e. size of the largest clique in $G$).
From Corollary 3.2, the average regret satisfies $R_{AB}(t)=\tilde{O}\left(\frac{\sqrt{\theta(G)\xi_{\omega(G)}\Psi_{t\omega(G)}}}{\sqrt{Mt}}\right)$ (recall $\omega(G)$ denotes the clique number of $G$). We note that the term $\Psi_{t\omega(G)}$ corresponds to the maximal mutual information gain from $t\omega(G)$ observations, and that this quantity depends only logarithmically on $t\omega(G)$ for standard kernels such as the squared exponential kernel. For more details, see Appendix B. The term $\xi_{\omega(G)}$ is the price we pay for the absence of coordination within each of the subgraphs $G_{k}$ in $G$, and is a standard term that arises in multi-agent Bayesian optimization. By an appropriate initialization phase, this term can be reduced to $\tilde{O}(1),$ (see Appendix B.3 in [24]). Thus, compared to the sequential single-agent case with $t$ rounds which has average regret $\tilde{O}\left(\sqrt{\frac{1}{t}}\right)$ [14], our algorithm satisfies a regret of $\tilde{O}\left(\frac{\sqrt{\theta(G)}}{\sqrt{Mt}}\right)$, i.e. an improvement of $\sqrt{\frac{\theta(G)}{M}}$ (note this term is always smaller than 1). Correspondingly, the average regret is smaller for graphs with higher connectivity, whose clique cover number $\theta(G)$ is smaller. We next proceed to bound the Bayesian simple regret.
Figure 1: Regret analysis of numerical simulations with $20$ agents on Erdős-Rényi random graphs with connectivity probability $0.2$ (blue), $0.4$ (orange), and $0.6$ (green), on Rosenbrock (left) and Ackley (right) objective functions.
@@
Our proof begins by observing that the Bayesian simple regret is bounded by the Bayesian average regret of any subset of agents. Similar to the proof for Theorem 3.1, we upper bound Bayesian average regret on a specific subset of agents by decomposing regret into three sums, each of which utilize the confidence function $U_{t,i}.$ The bounds follow notions from information theory using techniques from Kandasamy et al. [22], incorporating the information gain constants introduced in Equations 3 and 4.
@@
Recall from equation 2 that $R_{SB}(t)=\min_{i\in\{1,2,\ldots,M\},\tau\in\{1,2,\ldots,t\}}\mathbb{E}[f(x^{*})-f(x_{\tau,i})].$ Define $U_{t,i}(x)=\mu_{D_{t,i}}(x)+\beta_{t}^{1/2}\sigma_{D_{t,i}}(x).$ Note that, since $R_{SB}(t)$ represents a minimum value of the expression $\mathbb{E}[f(x^{*})-f(x_{\tau,i})],$ it is upper bounded by the average of this expression across any subset of agents. Thus, we may write
$R_{SB}(t)\leq\frac{1}{t|V_{s}|}\sum_{\tau=1}^{t}\sum_{i\in V_{s}}\mathbb{E}[f(x^{*})-f(x_{\tau,i})].$
Thus, it suffices to upper bound $\frac{1}{t|V_{s}|}\sum_{\tau=1}^{t}\sum_{i\in V_{s}}\mathbb{E}[f(x^{*})-f(x_{\tau,i})].$ Note that this expression is equal to the Bayesian average regret over $|V_{s}|$ agents on the complete subgraph $G_{s}=(V_{s},E_{s}).$ By Theorem 3.1, we can bound Bayesian average regret; thus, we may write $\frac{1}{t|V_{s}|}\sum_{\tau=1}^{t}\sum_{i\in V_{s}}\mathbb{E}[f(x^{*})-f(x_{\tau,i})]\leq\frac{\sqrt{2}\pi^{3/2}}{12t|V_{s}|}+\sqrt{\frac{2\xi_{|V_{s}|}\beta_{t}\Psi_{t|V_{s}|}}{t|V_{s}|\log(1+\sigma_{\epsilon}^{-2})}}.$ Because $R_{SB}(t)\leq\frac{1}{t|V_{s}|}\sum_{\tau=1}^{t}\sum_{i\in V_{s}}\mathbb{E}[f(x^{*})-f(x_{\tau,i})],$ $R_{SB}(t)$ inherits this bound. Therefore, we have shown $R_{SB}(t)\leq\frac{C_{1}}{t|V_{s}|}+\sqrt{\frac{C_{2}\xi_{|V_{s}|}\beta_{t}\Psi_{t|V_{s}|}}{t|V_{s}|}},$ where $\beta_{t}=2\log(t^{2}|V_{s}||\mathcal{X}|),$ $C_{1}=\frac{\sqrt{2}\pi^{3/2}}{12},$ and $C_{2}=\frac{2}{\log(1+\sigma_{\epsilon}^{-2})},$ concluding our proof.
@@
Picking $G_{s}$ to be the largest complete subgraph of the communication network $G$ then yields the following corollary.
@@
Suppose $k(x,x^{\prime})\leq 1$ for all $x,x^{\prime}.$ Let $G_{max}=(V_{max},E_{max})$ be the largest complete subgraph of $G.$ Then the Bayesian simple regret after $t$ timesteps satisfies $R_{SB}(t)\leq\frac{C_{1}}{t|V_{max}|}+\sqrt{\frac{C_{2}\xi_{|V_{max}|}\beta_{t}\Psi_{t|V_{max}|}}{t|V_{max}|}},$ where $\beta_{t}=2\log(t^{2}|V_{max}||\mathcal{X}|),$ $C_{1}=\frac{\sqrt{2}\pi^{3/2}}{12},$ and $C_{2}=\frac{2}{\log(1+\sigma_{\epsilon}^{-2})}.$
From the above corollary and our discussion following Corollary 3.2, we see that $R_{SB}(t)=\tilde{O}\left(\sqrt{\frac{1}{t|V_{max}|}}\right)$. Thus, compared to the sequential single-agent case with $t$ rounds which has simple regret $\tilde{O}\left(\sqrt{\frac{1}{t}}\right)$ [14], our algorithm satisfies a regret of $\tilde{O}\left(\sqrt{\frac{1}{t|V_{max}|}}\right)$, i.e. an improvement of $\sqrt{\frac{1}{|V_{max}|}}$, demonstrating the benefit of the network structure for the simple regret case as well.
IV Numerical Experiments
IV-A Simulation
In the numerical implementation, performance was assessed utilizing the following regret metrics. We define the Instant average regret $R_{A}$, and its sum, $\overline{R_{A}}$, as follows: